Can you help solve this system?
y=3^(2x+1)
y=9^(4-2x)
I know 9=3^2, so y=(3^2)^(4-2x)
Thank you.

Answer by stanbon(75887)

y=3^(2x+1)
y=9^(4-2x)
I know 9=3^2, so y=(3^2)^(4-2x) , so y = 3^(8-4x)
-------
Therefore::
3^(2x+1) = 3^(8-4x)
----
Then::
2x+1 = 8-4x
6x = 7
x = 7/6
-------------------
Cheers,
Stan H.